# Training Gradient Boosted Decision Trees on Tabular Data Containing Label Noise for Classification Tasks

### v1

@@ 5.2 Effect of Label Noise on Training GBDTs
This section deals with the first set of research questions, namely the effects of label noise on GBDTs without any methods correcting for label noise. To get a measure for the model’s classification performance, the performance of the trained model was evaluated on a clean test set, unpolluted by noise, and will be denoted as “test” in the following figures.
Both the Covertype and the Dry Bean dataset have more than two classes and were selected because it is generally easier to extrapolate findings from multiclass classification to binary classification than vice versa. The Covertype dataset is very large and complex, which is why it was chosen as the main dataset to generate the figures from. Using such datasets in research resembles real-world situations better, ensuring findings of this work aren’t just based on basic patterns but are more adaptable to varied scenarios.
However, sometimes the Covertype dataset was too large to calculate the metrics for all noise detection and treatment methods across all noise types and rates. In such cases, the Dry Bean dataset was used for these calculations instead. If there isn’t a significant difference in the plots generated for pair and symmetric noise, only one of them will be presented in this work.
To find out how naturally robust GBDTs are to label noise (RQ1) and what types of errors the model makes as training progresses (RQ2), the model was first trained without early stopping for 100 epochs and its predictions were collected and categorized into three groups: the prediction matches (1) the ground-truth label, (2) the noisy label or (3) another class altogether. Figure 5 shows that for both noise types, the model correctly predicts the true class most of the time at the start of training. The number of correct predictions decreases as training progresses and the number of predictions that match the noisy label grows. At about 60 epochs with 10% pair noise and 95 epochs with symmetric noise, predictions that match with the true and noisy label are equally likely. Only few predictions match a label that is neither the noisy nor the true label in the beginning and this number decreases as training progresses.
@@
From those two figures one can infer that GBDTs, despite the fact that boosting algorithms are known to be very sensitive to noise, actually mostly predict the correct label and only gradually adapt to the noisy labels (RQ1, RQ2). Therefore early stopping is an important tool to use for training.
@@
Figure 7 illustrates the model’s log loss per epoch, with each line representing training at a different noise rate. Lighter hues indicate lower noise rates, while darker hues represent higher ones. All training curves reach a minimum at approximately 8 epochs, with the darker lines converging at a significantly higher loss than the lighter ones. At 0% noise, the model achieves a similar log loss on both noise types. However, starting from noise rates as low as 10%, the final log loss when trained on pair noise is noticeably higher than when trained on symmetric noise. This indicates that model performance on the test set deteriorates as the noise rate increases and that pair noise presents a greater challenge for learning at noise rates higher than 10%. The reason behind this is that with pair noise, the label is flipped to only one class whereas with symmetric noise, the noise rate is equally distributed among the other classes. Therefore pair noise has a stronger influence when training the model.
Figure 8 plots the model’s accuracy against the noise rate it was trained on. For both symmetric and pair noise, the line has a downward slope, indicating that accuracy decreases as the noise rate increases. For pair noise, the line drops drastically at a noise rate of 50%, which is to be expected since from 50% pair noise upwards, predictions are random or misclassification more likely than the true class. However, prior to that point, the lines for symmetric and pair noise are quite similar, suggesting that training on either type of noise has a comparable impact on the model’s accuracy when early stopping is used.
Both Figure 7 and Figure 8 highlight the grave effect of label noise on classifier performance and the importance of investigating methods to counteract it.
An interesting point is that the slower decrease in test set accuracy when training on symmetric rather than pair noise (Figure 6) indicates that GBDTs are more robust to symmetric than pair noise.
Figure 7: Logloss on the test set per epoch at varying noise rates (Dry Bean). Logloss increases as the noise rate increases.
Figure 8: Classification accuracy on the test set at different noise rates (Dry Bean). Classification accuracy decreases as the noise rate increases.
Figure 9 shows the training curves when training on 30% symmetric and pair noise, respectively. With symmetric noise, performance on the training set converges earlier and at a higher logloss than with pair noise, indicating that symmetric noise is harder to learn with. On the other hand, final performance on the test set is in turn slightly higher with symmetric noise than with pair noise, signifying that the model would still perform well if the source of the label noise were to be removed or a label noise treatment method were to be employed.
Figure 9: Training curves at 30% noise (Covertype). Final test set performance is slightly higher when trained on symmetric noise.
@@
In summary, while symmetric noise appears harder for the model to learn, it results in better logloss performance on a clean test set (RQ3). The training curves (Figure 6) confirm that, depending on when training is stopped, training on symmetric noise results in a higher classification accuracy on the test set than training on pair noise.
5.3 Performance Comparison
This section deals with the second set of research questions, which aim to compare the performance of the implemented noise detection and correction methods.
@@ 5.3 Performance Comparison
From this, one might infer that AUM and Gradients could be the most effective in separating clean from noisy instances (RQ4), given their unimodal distributions. Their performance might be followed by ConfCorr, whose distributions are bimodal and have more overlap but still maintain a considerable gap. Although LRT clusters all clean instances closely around a single value, the presence of some noisy instances within this distribution suggests that LRT’s performance might be on the lower end.
@@
To further investigate which methods detect noise best (RQ4), one can examine the confusion matrices comparing actual versus predicted noisy instances (Figure 12). All methods accurately identified noisy and clean instances approximately 80% of the time. Gradients achieved the highest performance, as it has the highest sum on the diagonal of the confusion matrix.
Figure 12: Label noise detection confusion matrices of each method on 30% pair noise (Covertype). Rows represent the ground truth and columns the predictions. Clean instances are represented by zero and noisy instances by one. All methods accurately identify noisy and clean instances approximately 80% of the time.
Figure 13 shows how the mean of the calculated metrics changes depending on the noise rate. AUM’s, ConfCorr’s and Gradients’ metric both converge as the noise rate increases, demonstrating that clean and noisy instances become more difficult to distinguish as the noise rate increases (RQ5). All metrics intersect and then diverge at a noise rate of 50% and beyond, which was expected since misclassification becomes as or more likely than the true class. The values calculated for clean instances remain close to zero with LRT and those for noisy instances approach zero as the noise rate increases. In the case of symmetric noise, the lines would intersect at a noise rate of $\frac{c-1}{c}$.
The convergence and eventual intersection of these lines demonstrate that for all methods, differentiating between clean and noisy instances becomes increasingly difficult as the noise rate approaches 50% (RQ5).
Figure 13: Mean metrics calculated at different noise rates with pair noise (Dry Bean). Most methods converge as the noise rate increases, meaning that clean and noisy instances become increasingly harder to differentiate.
@@
To summarize, LRT and AUM achieve the highest noise detection accuracy across all noise rates on the Dry Bean dataset, closely followed by Gradients and ConfCorr (RQ4).
@@
Figure 15 maps the actual noise rate against the percentage of instances marked as noisy by AUM at the first epoch after warm-up. The black dotted line represents an ideal noise detection classifier, where the percentage of instances marked as noisy matches the actual noise rate present in the data. On the Dry Dean dataset, a few instances are marked as noisy even though no noise is present. For both pair and symmetric noise, the line stays very close to the “ideal” black dotted line. From a noise rate of 40% onward, AUM predicts about 35% of the instances to be noisy even as the noise rate increases. On the Covertype dataset, the two lines generally lie above the ideal, meaning that usually more instances are marked as noisy than are actually present in the dataset.
Therefore on simpler datasets with few outliers and little class imbalance, the percentage of instances marked as noisy may be used to approximate the amount of noise present in the data, especially if the noise is symmetric (RQ5).
Figure 15: Noise detection accuracy of AUM at various noise rates, calculated at the first epoch after warm-up. AUM marks about as many instances as noisy as are actually present in on the Dry Bean dataset, but tends to mark more as noisy as are actually present on the Covertype dataset. Therefore the proportion of instances marked as noisy can be used as an estimate for the noise rate on simpler datasets.
Similarly, Figure 16 shows the detection accuracy of all methods on the Dry Bean dataset at the first epoch after warm-up. LRT and AUM overlap again, which is why AUM’s line isn’t visible.
The most prominent difference to the description above is that Gradients’ percentage of marked instances drops rapidly at a pair noise level of 50%. Both Gradients and ConfCorr tend to label more instances as noisy than are actually present in the datasets. This could indicate that they are more sensitive to outliers or class imbalances. In contrast, LRT and AUM generally mark a number of instances as noisy that is close to the actual amount present in the data. Consequently, the percentage of instances marked as noisy by LRT and AUM can be used to approximate the level of noise present in simple dataset polluted by symmetric noise (RQ5). For pair noise exceeding 40%, these estimates become increasingly unreliable.
@@
This means that metrics for noisy and clean instances tend to converge as training progresses (RQ6). Depending on the metric, both values either increase or decrease. The values calculated on noisy instances tend to have a steeper slope, meaning that they change more drastically during training than those calculated on clean instances.
The ability of all methods to discriminate between clean and noisy instances remains roughly constant during training, with the exception of LRT. The convergence for LRT’s plot shows that the model adapts to noisy instances as training progresses and therefore noisy and clean instances become increasingly harder to differentiate (RQ6), highlighting again that early stopping is an important tool to avoid overfitting noisy labels. For all methods, using a fixed threshold without early stopping will likely be ineffective in discriminating between clean and noise instances throughout the whole training procedure.
@@
Removal and relabeling appear to perform roughly the same according to Figure 18. The drop in accuracy is slightly lower with relabeling as well as the increase in error after about twenty epochs. Based solely on this information, a definitive conclusion regarding the best correction method (RQ7) cannot be determined. A numeric instead of a visual comparison of all noise detection and correction methods in Section 5.4 will give a clearer perspective on how to definitively answer which noise correction method performs better in which situations.
@@
As mentioned previously, a numeric comparison in the next subsection will clarify whether removing or relabeling suspicious instances reaches a better classification performance.
@@
5.4 Final Results
A visual comparison between the implemented noise detection and correction methods can give a quick impression of their differences, but numeric data is needed for a more detailed and nuanced comparison. This section therefore presents the noise detection and classification performance results in various noise settings and on different datasets. Here, the performance of the implemented methods will also be compared to state-of-the-art research using the same metrics, datasets and similar noise settings. The following tables were generated at the most optimal epoch as indicated by early stopping. In them, the best values per category are marked in bold letters.
In their paper, Ponti et al. (2022) implemented a variant of pair noise. On the Adult dataset, they achieved a maximum noise detection accuracy of 86.93%, 83.97% and 82.58% at noise rates from 10% to 30% each. On the same dataset, the methods implemented in this work achieve a maximum noise detection accuracy of 99.70%, 99.20% and 98.56% at noise rates from 10% to 30% each (Table 3), thereby outperforming Ponti et al.’s work across all noise rates.
This work’s noise detection accuracy on the Covertype dataset on noise rates from 10% to 30% are 83.06%, 82.90% and 81.00% each (Table 4) and is thereby outperformed by Ponti et al’s work which reached a noise detection accuracy of 91.15%, 89.12% and 85.95%.
On the Breast Cancer dataset, Ponti et al. achieved a maximum noise detection accuracy of 95.57%, 93.68% and 89.68% at noise rates from 10% to 30% each. This work, though, underperforms on the Breast Cancer dataset, achieving a maximum noise detection accuracy of 92.06%, 90.75% and 85.25% at the three noise levels (Table 5).
The performance difference between this paper and Ponti et al.’s work when using ConfCorr may partially be due to the difference in the pair noise implementation mentioned above and in the way the threshold is chosen.
@@ 5.4 Final Results
Noise Rate Detection Accuracy FNR FPR
0.10 AUM 99.70 0.29 0.29
ConfCorr 93.30 6.70 6.70
Gradients 94.00 6.02 6.02
LRT 99.70 0.29 0.29
0.20 AUM 99.20 0.80 0.80
ConfCorr 94.40 5.60 5.60
Gradients 98.00 2.03 2.03
LRT 99.20 0.80 0.80
0.30 AUM 98.50 1.48 1.48
ConfCorr 98.44 1.58 1.58
Gradients 98.56 1.46 1.46
LRT 98.50 1.48 1.48
@@
Noise Rate Detection Accuracy FNR FPR
0.10 AUM 80.94 19.10 19.10
ConfCorr 83.06 16.94 16.94
Gradients 77.20 22.80 22.80
LRT 80.94 19.10 19.10
0.20 AUM 80.90 19.16 19.16
ConfCorr 82.90 17.16 17.16
Gradients 79.90 20.11 20.11
LRT 80.90 19.16 19.16
0.30 AUM 79.60 20.34 20.34
ConfCorr 81.00 19.00 19.00
Gradients 78.80 21.16 21.16
LRT 79.60 20.34 20.34
@@
Noise Rate Detection Accuracy FNR FPR
0.10 AUM 90.30 9.67 9.67
ConfCorr 92.06 7.91 7.91
Gradients 90.56 9.45 9.45
LRT 90.30 9.67 9.67
0.20 AUM 80.44 19.56 19.56
ConfCorr 90.75 9.23 9.23
Gradients 90.75 9.23 9.23
LRT 80.44 19.56 19.56
0.30 AUM 72.06 27.90 27.90
ConfCorr 85.25 14.73 14.73
Gradients 82.60 17.36 17.36
LRT 72.06 27.90 27.90
@@
Zheng et al. (2020) applied 10% pair noise in a way that doesn’t affect the class distribution, so the noise implementation again differs a little from this work. They tested their work on multiple small-sized datasets, from which one is used in this work as well. On the Breast Cancer dataset, they achieved a classification accuracy of 94.9%, outperforming this work by about 1.9% (Table 8).
@@
Noise Rate Noise Type Correction Detection Accuracy Precision Recall F1
0.10 pair none none 90.40 96.94 87.50 92.00
@@
0.30 none none 80.70 91.70 76.40 83.30
relabel AUM 79.80 90.20 76.40 82.70
ConfCorr 79.80 90.20 76.40 82.70
Gradients 79.80 90.20 76.40 82.70
LRT 79.80 90.20 76.40 82.70
remove AUM 78.94 91.40 73.60 81.56
ConfCorr 89.50 94.10 88.90 91.44
Gradients 89.50 92.90 90.25 91.56
LRT 78.94 91.40 73.60 81.56
@@
Table 8: Classification performance on the Breast Cancer dataset with 10% and 30% pair noise. Highest values are marked by noise rate.
Table 9 displays the classification performance of all methods at different noise rates and types. On 20% pair noise, Gradients together with noisy candidate removal achieved the highest classification accuracy, precision, recall and f-1 score. On 20% symmetric noise, AUM combined with relabeling achieved the highest classification accuracy, but ConfCorr combined with removal the highest precision, recall (along with Gradients) and f1-score. On 40% pair noise, Gradients and noisy candidate removal reached the highest performance across all metrics (some on par with ConfCorr), but on symmetric noise, AUM combined with relabeling reached the highest classification performance metrics.
These results indicate that the most effective combination of noise detection and correction methods depends on the specific dataset, the type and the amount of noise present in the data.
Noise Rate Noise Type Correction Detection Accuracy Precision Recall F1
0.20 pair none none 90.60 91.40 90.56 90.94
relabel AUM 91.56 93.00 91.56 92.25
ConfCorr 91.70 93.06 91.80 92.40
Gradients 91.60 93.06 91.80 92.40
LRT 91.56 93.00 91.56 92.25
remove AUM 91.56 93.00 91.56 92.25
ConfCorr 92.00 93.60 93.06 93.30
Gradients 92.10 93.75 93.20 93.50
LRT 91.56 93.00 91.56 92.25
symmetric none none 91.50 93.00 92.10 92.56
relabel AUM 91.90 93.44 92.75 93.06
ConfCorr 91.75 93.44 92.70 93.00
Gradients 91.75 93.44 92.70 93.00
LRT 91.75 93.40 92.56 92.94
remove AUM 91.75 93.30 92.60 92.94
ConfCorr 91.75 93.50 92.90 93.20
Gradients 91.75 93.40 92.90 93.10
LRT 91.75 93.30 92.60 92.94
0.40 pair none none 78.25 75.50 76.56 75.94
relabel AUM 86.94 86.80 85.70 86.10
ConfCorr 86.90 86.44 85.70 86.00
Gradients 86.70 86.00 85.40 85.60
LRT 87.00 86.75 85.75 86.10
remove AUM 87.30 87.06 86.00 86.44
ConfCorr 88.30 88.44 87.00 87.56
Gradients 88.60 88.44 87.10 87.70
LRT 87.25 87.06 86.06 86.40
symmetric none none 91.00 92.25 91.56 91.90
relabel AUM 91.90 93.44 92.75 93.06
ConfCorr 91.56 93.06 92.40 92.70
Gradients 91.56 93.06 92.30 92.70
LRT 91.70 93.25 92.56 92.90
remove AUM 91.90 93.25 92.56 92.90
ConfCorr 91.20 92.56 92.00 92.30
Gradients 91.25 92.70 92.10 92.40
LRT 91.75 93.25 92.44 92.80
Table 9: Classification performance on the Dry Bean dataset with noise rates 20% and 40%. Highest values are marked by noise type.
6 Results and Discussion
This section summarizes the experimental findings, their implications, and evaluates the validity of the results in relation to the research questions posed in Section 2.5. The experiments showed that GBDTs are robust to label noise, especially symmetric noise, reaching a local minimum on the clean test set quickly, suggesting that early stopping can yield a good model.
The methods AUM and LRT achieved a noise detection accuracy of 99.30% on the Adult dataset. Differentiating noisy and clean instances becomes harder as noise increases. The metrics for clean and noisy instances change during training, so a dynamic threshold or early stopping is advisable. The effectiveness of noise detection and correction methods varies by dataset and noise type.
The Covertype dataset, with its multiple classes and high class imbalance, performed differently than the binary Adult dataset. Although LRT struggled with separating clean from noisy instances, it consistently assigned low values to clean instances, indicating effectiveness. AUM achieved the highest classification accuracy on symmetric noise on the Dry Bean dataset when paired with relabeling.
ConfCorr effectively separates clean from noisy instances, though its optimal threshold depends on noise rate, type, and dataset. A two-component GMM can work around this but is less efficient. ConfCorr and Gradients attained high classification precision on the Adult dataset and recall on the Breast Cancer dataset.
Though Gradients initially seemed less promising for noise detection, it achieved high precision and recall on different datasets. Similar to ConfCorr, it faces challenges with determining an optimal threshold.
In summary, the choice of detection and correction methods is influenced by noise rate, type, and dataset. Class imbalance may render some methods unsuitable, as they might remove instances from smaller classes. Relabeling can sometimes introduce more errors, while removing suspicious instances might discard informative data. It’s generally advisable to treat fewer instances than the estimated amount of noise present in the data. Multiple methods should be tested to find the optimal solution for label noise problems.
Due to time and computational constraints, each experiment was performed only once. The variability in method performance across different datasets suggests that favorable outcomes might be due to chance rather than inherent effectiveness. Cross-validation could address these concerns. Additionally, pair and symmetric noise may not represent real-world label noise, limiting the applicability of these findings.

### v2

@@ 6.1 Effect of Label Noise on Training GBDTs
This section examines the effects of label noise on GBDTs when not correcting for label noise. To get a measure for the model’s classification performance, the performance of the trained model was evaluated on a clean test set, unpolluted by noise, and is denoted as “test” in the following figures.
The Covertype and Dry Bean datasets, both multiclass, were chosen to enable generalization to binary classification tasks. The complex Covertype dataset served as the primary source for figures, with the Dry Bean dataset used for cases where Covertype’s size proved computationally limiting. Only one representative noise type (symmetric or pair) is shown if both produced similar trends.
To find out how naturally robust GBDTs are to label noise and what types of errors the model makes as training progresses, the model was first trained without early stopping for 100 epochs and its predictions were collected and categorized into three groups: the prediction matches (1) the ground-truth label, (2) the noisy label or (3) another class altogether. Figure 3 shows that for both noise types, the model correctly predicts the true class most of the time at the start of training. The number of correct predictions decreases as training progresses and the number of predictions that match the noisy label grows. At about 60 epochs with 10% pair noise and 95 epochs with symmetric noise, predictions that match with the true and noisy label are equally likely. Only few predictions match a label that is neither the noisy nor the true label in the beginning and this number decreases as training progresses.
@@
From those two figures one can infer that GBDTs, despite the fact that boosting algorithms are known to be very sensitive to noise, actually mostly predict the correct label and only gradually adapt to the noisy labels. Therefore early stopping is an important tool to use for training.
@@
In summary, while symmetric noise appears harder for the model to learn, it results in better logloss performance on a clean test set. The training curves (Figure 4) confirm that, depending on when training is stopped, training on symmetric noise results in a higher classification accuracy on the test set than training on pair noise.
6.2 Noise Detection and Correction
In this section, the performance of the implemented noise detection and correction methods is compared.
@@ 6.2 Noise Detection and Correction
From this, one might infer that AUM and Gradients could be the most effective in separating clean from noisy instances, given their unimodal distributions. Their performance might be followed by ConfCorr, whose distributions are bimodal and have more overlap but still maintain a considerable gap. Although LRT clusters all clean instances closely around a single value, the presence of some noisy instances within this distribution suggests that LRT’s performance might be on the lower end.
@@
To summarize, LRT and AUM achieve the highest noise detection accuracy across all noise rates on the Dry Bean dataset, closely followed by Gradients and ConfCorr.
@@
Figure 8 maps the actual noise rate against the percentage of instances marked as noisy by the noise detection methods at the first epoch after warm-up. The black dotted line represents an ideal noise detection classifier, where the percentage of instances marked as noisy matches the actual noise rate present in the data.LRT and AUM overlap, which is why AUM’s line isn’t visible.
LRT and AUM generally mark a number of instances as noisy that is close to the actual amount present in the data. At the lower noise rates, both Gradients and ConfCorr tend to label more instances as noisy than are actually present in the datasets. This could indicate that they are more sensitive to outliers or class imbalances. Consequently, the percentage of instances marked as noisy by LRT and AUM can be used to approximate the level of noise present in datasets polluted by symmetric noise. For pair noise exceeding 40%, these estimates become increasingly unreliable.
@@
This means that metrics for noisy and clean instances tend to converge as training progresses. Depending on the metric, both values either increase or decrease. The values calculated on noisy instances tend to have a steeper slope, meaning that they change more drastically during training than those calculated on clean instances.
The ability of all methods to discriminate between clean and noisy instances remains roughly constant during training, with the exception of LRT. The convergence for LRT’s plot shows that the model adapts to noisy instances as training progresses and therefore noisy and clean instances become increasingly harder to differentiate, highlighting again that early stopping is an important tool to avoid overfitting noisy labels. For all methods, using a fixed threshold without early stopping will likely be ineffective in discriminating between clean and noise instances throughout the whole training procedure.
@@
Removal and relabeling appear to perform roughly the same according to Figure 10. The drop in accuracy is slightly lower with relabeling as well as the increase in error after about twenty epochs. Based solely on this information, a definitive conclusion regarding the best correction method cannot be determined.
@@
6.3 Comparison with the State-of-the-Arts
This section numerically presents the noise detection and classification performance results in various noise settings and on different datasets. The performance of the implemented methods is also compared to state-of-the-art research using the same metrics, datasets and similar noise settings. The following tables were generated at the most optimal epoch as indicated by early stopping. In them, the best values per category are marked in bold letters.
In their paper, Ponti et al. (2022) implemented a variant of pair noise. On the Adult dataset, they achieved a maximum noise detection accuracy of 86.93%, 83.97% and 82.58% at noise rates from 10% to 30% each. On the same dataset, the methods implemented in this work achieve a maximum noise detection accuracy of 99.70%, 99.20% and 98.56% at noise rates from 10% to 30% each (Table 3), thereby outperforming Ponti et al.’s work across all noise rates. But none of the implemented methods managed to outperform their work on the Covertype and Breast Cancer datasets.
@@ 6.3 Comparison with the State-of-the-Arts
Noise Rate Detection Accuracy
0.10 AUM 99.70
ConfCorr 93.30
Gradients 94.00
LRT 99.70
0.20 AUM 99.20
ConfCorr 94.40
Gradients 98.00
LRT 99.20
0.30 AUM 98.50
ConfCorr 98.44
Gradients 98.56
LRT 98.50
@@
Noise Rate Detection Accuracy
0.10 AUM 80.94
ConfCorr 83.06
Gradients 77.20
LRT 80.94
0.20 AUM 80.90
ConfCorr 82.90
Gradients 79.90
LRT 80.90
0.30 AUM 79.60
ConfCorr 81.00
Gradients 78.80
LRT 79.60
@@
Noise Rate Detection Accuracy
0.10 AUM 90.30
ConfCorr 92.06
Gradients 90.56
LRT 90.30
0.20 AUM 80.44
ConfCorr 90.75
Gradients 90.75
LRT 80.44
0.30 AUM 72.06
ConfCorr 85.25
Gradients 82.60
LRT 72.06
@@
Correction Detection Accuracy Precision Recall F1
none none 90.40 96.94 87.50 92.00
@@
Table 7: Classification performance on the Breast Cancer dataset with 30% pair noise.
6.4 Summary
Our results demonstrate that GBDTs show a natural robustness to label noise, particularly to symmetric noise. By implementing early stopping, we effectively prevent GBDTs from adapting to noisy labels, thereby preserving test accuracy.
Both AUM and LRT achieved state-of-the-art noise detection accuracy, reaching 99% across all noise rates on the Adult dataset, and proved effective for estimating the amount of noise present in the data. In addition, ConfCorr demonstrated high classification precision on the Adult dataset, while Gradients achieved superior classification precision and recall on the Covertype and Breast Cancer datasets, respectively, outperforming other approaches.
The use of dynamic thresholding with a Gaussian Mixture Model (GMM) for ConfCorr and Gradients further enhanced performance by eliminating the need for additional parameter tuning, allowing these methods to adapt to varying noise levels automatically.
7 Discussion
Our findings suggest that noise detection performance is highly dependent on dataset characteristics, noise type, and noise rate, indicating that the optimal combination of detection and correction methods is context-specific. Class imbalance also presents challenges, as some methods may inadvertently remove instances from smaller classes, potentially reducing model performance.
Relabeling noisy instances can sometimes introduce additional errors, while removing them might lead to the loss of informative data. Consequently, multiple methods should be evaluated to identify the best approach for managing label noise in any given scenario. A cautious approach is recommended, generally treating fewer instances than the estimated noise level to mitigate potential drawbacks.
All methods exhibited sensitivity to class imbalance, with performance notably lower on the highly imbalanced Covertype dataset. Furthermore, time and computational constraints limited the study to a single trial per experiment. The observed variability in method performance across datasets suggests that favorable outcomes may occasionally result from chance rather than inherent effectiveness, underscoring the value of cross-validation for greater reliability.
Finally, the use of simple noise types, such as pair and symmetric noise, may limit the real-world applicability of these results, as they do not fully represent the complexity of real-world label noise.